OpenStudy (anonymous):

Find the absolute minimum of the function of f(x) = 3 x^2 - 10 x + 6, in the interval 0 is less than or equal to x and x is less than or equal to 8.

OpenStudy (zzr0ck3r):

can you find the derivative?

OpenStudy (anonymous):

\[f \prime(x)=6x-10\]

OpenStudy (zzr0ck3r):

set that equal to 0 and solve for x

OpenStudy (zzr0ck3r):

this is a upward opening parabola, so the min is the vertex...

OpenStudy (anonymous):

x= -10/6

OpenStudy (zzr0ck3r):

why -?

OpenStudy (anonymous):

i meant 10/6 whooopsie

OpenStudy (zzr0ck3r):

yep

OpenStudy (anonymous):

would that be the min?

OpenStudy (zzr0ck3r):

do you know how to find the vertex of a parabola?

OpenStudy (zzr0ck3r):

-b/(2a)

OpenStudy (zzr0ck3r):

yes its a min

OpenStudy (anonymous):

that ended up being wrong.

OpenStudy (zzr0ck3r):

its not, its the only min....

OpenStudy (zzr0ck3r):

its the vertex

OpenStudy (zzr0ck3r):

-b/(2a) = 10/6

OpenStudy (anonymous):

huh? so how do I find the absolute min?

OpenStudy (zzr0ck3r):

same thing

OpenStudy (zzr0ck3r):

think of what a parabola looks like

OpenStudy (zzr0ck3r):

its smallest value is the vertex right?

OpenStudy (zzr0ck3r):

this is also where it has a slope = 0

OpenStudy (zzr0ck3r):

so there are two ways to find the max/min of a parabola

OpenStudy (rsadhvika):

put x = 10/6 in f(x) = 3x^2-10x+6

OpenStudy (zzr0ck3r):

yes you found where its min is, you need to plug that back into f(x) to get the value

OpenStudy (zzr0ck3r):

you found the x coordinate...

OpenStudy (anonymous):

i got -2.3333333 and it ended up wrong again.. @zzr0ck3r

OpenStudy (zzr0ck3r):

don't round

OpenStudy (zzr0ck3r):

-7/3

OpenStudy (anonymous):

ohhhh okay. I got it!!! (: thank you so much! @zzr0ck3r

OpenStudy (zzr0ck3r):

np